### v1

@@ 3.2.1 Prior distribution for $\mathcal{S}$
Given that $\mathcal{S}$ represents a sparse variable, we adopt a Laplace prior distribution to model its sparsity. The Laplace prior heavily penalizes large values of $\mathcal{S}$ while allowing for many small deviations from zero, thereby promoting sparsity in the estimated solution. The Laplace prior is parameterized by $\theta_{2}$ and is formally defined as:
(8) $p(\mathcal{S}|\theta_{2})\propto{\theta_{2}^{n}}\exp\left(-\theta_{2}\|\mathcal{S}\|_{1}\right).$
This characteristic encourages sparsity in $\mathcal{S}$ and is in alignment with the regularization term $\left\|\mathcal{S}\right\|_{1}$ in (2). Taking the logarithm of the prior distribution, we obtain:
(9) $\log p(\mathcal{S}|\theta_{2})=-\theta_{2}\|\mathcal{S}\|_{1}+n\log\theta_{2}+C_{2},$
@@ 3.2.2 Prior distribution for $\mathcal{L}$
For the variable $\mathcal{L}$, we employ a Gibbs prior [28], which promotes a low-rank structure in $\mathcal{L}$. The Gibbs prior, also known as an exponential prior with a tensor nuclear norm penalty, is a convex surrogate for the tensor rank. It encourages $\mathcal{L}$ to have a low-rank representation, which is often suitable for capturing the underlying low-dimensionality in the data. The prior distribution is given by:
(10) $p(\mathcal{L}|\theta_{3})\propto{\theta_{3}^{n}}\exp\left(-\theta_{3}\left\|\mathcal{L}\right\|_{*}\right).$
@@ 3.3 The hyper-prior distribution
In the field of statistical modeling, the Gamma distribution has obtained significant attention as a versatile prior distribution for hyperparameters, particularly in Bayesian frameworks [4, 2, 37, 3, 41]. This preference stems from its ability to represent non-negative values with flexible shape and scale parameters. The Gamma density function, given by
(11) $p(x)=\mathcal{G}(x|a,b)\propto x^{a-1}\exp(-bx),$
where $a>0$ and $b>0$ represent shape and scale parameters respectively. The mean and variance of these Gamma distributions are given by
(12) ${\mathbb{E}(x)}=\frac{a}{b},\quad{\mathbb{V}\mathrm{ar}(x)}=\frac{a}{b^{2}}.$
For the hyperparameters $\theta_{i}$, we assume independent Gamma priors:
@@ 3.4 Joint distribution
We remark that, in the MAP framework, the hyperparameters $\boldsymbol{\theta}=(\theta_{1},\theta_{2},\theta_{3})$ must be either pre-specified or estimated prior to the estimation of $\mathcal{L}$ and $\mathcal{S}$. For a more comprehensive estimation that includes the hyperparameters, the joint maximum a posteriori (JMAP) estimation is employed:
$(\mathcal{S}^{\dagger},\mathcal{L}^{\dagger},\boldsymbol{\theta}^{\dagger})=\arg\max_{\mathcal{S},\mathcal{L},\boldsymbol{\theta}}p(\mathcal{S},\mathcal{L},\boldsymbol{\theta}|\mathcal{X})=\mathop{\rm argmax}_{\mathcal{S},\mathcal{L},\boldsymbol{\theta}}\frac{p(\mathcal{X},\mathcal{S},\mathcal{L},\boldsymbol{\theta})}{p(\mathcal{X})}.$
For simplicity, we assume independence among the hyperparameters, allowing us to express the joint density function of the variables $\mathcal{X}$, $\mathcal{L}$, $\mathcal{S}$, and $\theta$ as:
@@
In this paper, we assign zero values to both the shape and rate parameters of the Gamma prior for the hyperparameters:
(13) $a_{\theta_{i}}=b_{\theta_{i}}=0,\quad i=1,2,3.$
This choice results in an improper non-informative prior, where the prior density $p(\theta_{i})$ becomes proportional to a constant (i.e., $p(\theta_{i})\propto 1$ for all $\theta_{i}$), reflecting a complete lack of informative prior beliefs. Such priors are commonly referred to as vague priors or weakly informative priors in statistical literature [14]. Hence we have
(14) $p(\mathcal{X},\mathcal{S},\mathcal{L},\boldsymbol{\theta})\propto\theta_{1}^{n/2}\theta_{2}^{n}\theta_{3}^{n}\exp\left(-\frac{\theta_{1}}{2}\left\|\mathcal{X}-\mathcal{S}-\mathcal{L}\right\|_{F}^{2}-\theta_{2}\left\|\mathcal{S}\right\|_{1}-\theta_{3}\left\|\mathcal{L}\right\|_{*}\right).$
@@ 4 Variational Bayesian inference
In Bayesian modeling, inference involves conditioning on observed data $\mathcal{X}$ and estimating the posterior density $p(\mathcal{L},\mathcal{S},\boldsymbol{\theta}|\mathcal{X})$. This task can be tackled via Markov Chain Monte Carlo (MCMC) sampling or optimization approaches. However, in this paper, we adopt variational inference as the methodological framework to approximate the latent variables $\mathcal{L}$ and $\mathcal{S}$, along with the parameter vector $\boldsymbol{\theta}$.
@@ 4.1 Kullback-Leibler divergence and evidence lower bound
The central goal of variational inference is to identify an optimal variational density $q(\mathcal{L},\mathcal{S},\boldsymbol{\theta})$ that closely approximates the posterior density $p(\mathcal{S},\mathcal{L},\boldsymbol{\theta}|\mathcal{X})$, thereby facilitating efficient inference on the latent variables and parameters [7].
Within this framework, we define a family of densities $\mathcal{Q}$ over the latent variables and parameters. Each candidate $q(\mathcal{S},\mathcal{L},\boldsymbol{\theta})\in\mathcal{Q}$ represents an approximation to the true posterior. Let $\mathcal{Q}_{\mathcal{G}}$ represent the set of Gamma densities and $\mathcal{Q}_{\mathcal{N}}$ denote the set of normal densities. The overall variational family $\mathcal{Q}$ can be expressed as the Cartesian product of these sets: $\mathcal{Q}=\mathcal{Q}_{\mathcal{N}}\times\mathcal{Q}_{\mathcal{N}}\times\mathcal{Q}_{\mathcal{G}}$.
The optimal candidate is chosen by minimizing the Kullback-Leibler (KL) divergence from the true posterior:
[…]
@@
The posterior density $p(\mathcal{S},\mathcal{L},\boldsymbol{\theta}|\mathcal{X})$ is expressed as
$p(\mathcal{S},\mathcal{L},\boldsymbol{\theta}|\mathcal{X})=\frac{p(\mathcal{X},\mathcal{S},\mathcal{L},\boldsymbol{\theta})}{p(\mathcal{X})}.$
The density $p(\mathcal{X})$ involves integrating out the latent variables from the joint density. Unfortunately, this integration is often intractable, rendering direct computation of the posterior challenging. Expand the condition density, we have
@@
To circumvent the intractability, we optimize an alternative objective that is equivalent to the KL divergence up to an additive constant. Specifically, we minimize the first term on the right-hand side of the equation, which constitutes the evidence lower bound (ELBO), denoted $\mathcal{J}(q(\mathcal{S},\mathcal{L},\boldsymbol{\theta}))$
(16) $\mathcal{J}(q(\mathcal{S},\mathcal{L},\boldsymbol{\theta}))\equiv\mathbb{E}_{q(\mathcal{S},\mathcal{L},\boldsymbol{\theta})}\left[\log\left(\frac{p(\mathcal{X},\mathcal{S},\mathcal{L},\boldsymbol{\theta})}{q(\mathcal{S},\mathcal{L},\boldsymbol{\theta})}\right)\right].$
@@
(17) $q^{\dagger}(\mathcal{S},\mathcal{L},\boldsymbol{\theta})=\mathop{\rm argmax}_{{q(\mathcal{S},\mathcal{L},\boldsymbol{\theta})\in\mathcal{Q}}}\mathcal{J}(q(\mathcal{S},\mathcal{L},\boldsymbol{\theta})).$
@@ 4.2 Mean-field variational family
The selection of variational densities $q(\mathcal{L})$, $q(\mathcal{S})$, and $q(\theta_{i})$ is importance. For $q(\mathcal{L})$ and $q(\mathcal{S})$, we adopt normal distributions due to their versatility and analytical convenience. Insights from the likelihood function (7) and the priors specified in (8) and (10) indicate that the posterior distribution of $\theta_{1}$ and the conditional posteriors of $\theta_{i}$ for $i=2,3$ are Gamma-distributed. Thus, $q(\theta_{i})$ is naturally chosen as a Gamma density, ensuring compatibility with the posterior.
@@ 4.3 Laplacian approximation
In (14), the non-quadratic properties inherent in both the $\ell_{1}$ norm of $\mathcal{S}$, which represents the sum of the absolute values of all elements and the tensor nuclear norm of $\mathcal{L}$, which is the weighted sum of its singular values, pose significant obstacles for direct optimization within standard density families. These non-quadraticities complicate direct inference procedures, rendering them computationally intractable. To address this, we utilize the Laplace approximation method, involving mean calculation, variance estimation, and density function construction, to approximate the density with a Gaussian distribution form.
@@
where $\mathcal{Q}_{\mathcal{N}}$ is the set of all the density functions for Gaussian distribution. According to Gibbs’ inequality, for any two probability distributions $q(x)$ and $p(x)$ over a domain $\Omega$, the following holds:
@@ 4.3.1 Absolute value function
In the paper, we utilize the sans serif font $\mathsf{E}$ accompanied by a subscripted variable $x$ to denote the expectation of a variable $x$. The solution is given by:
@@
(18) $q(x)=\mathcal{N}\left(\mathsf{E}_{x},\frac{|\mathsf{E}_{x}|}{|\mathsf{E}_{x}|+\beta}\right).$
@@
(19) $\mathsf{E}|x|=|\mathsf{E}_{x}|+\frac{1}{2(|\mathsf{E}_{x}|+\beta)}.$
@@ 4.3.2 Nuclear norm
$\mathsf{E}_{\mathbf{X}}=\arg\min_{X}\left(\frac{\alpha}{2}\left\|\mathbf{X}-\mathbf{A}\right\|_{F}^{2}+\beta\left\|\mathbf{X}\right\|_{*}\right).$
@@
(20) $\left\|\mathbf{X}\right\|_{*}\approx\frac{1}{2}\text{Tr}(\omega(\mathsf{E}_{\mathbf{X}})\mathbf{X}\mathbf{X}^{T})+\frac{1}{2}\|\mathsf{E}_{\mathbf{X}}\|_{*}.$
@@
(21) $\boldsymbol{\Sigma}_{\mathbf{X}_{:j}}={\mathbf{U}}{\mathbf{D}}\left(\alpha{\mathbf{D}}+\beta\mathbf{I}\right)^{-1}{\mathbf{U}}^{T}.$
@@
(22) $q(\mathbf{X})=\prod_{j}\mathcal{N}(\mathbf{X}_{:j}|\mathsf{E}_{\mathbf{X}_{:j}},{\mathbf{U}}{\mathbf{D}}\left(\alpha{\mathbf{D}}+\beta\mathbf{I}\right)^{-1}{\mathbf{U}}^{T}).$
@@
(23) $\mathbb{E}\left\|\mathbf{X}\right\|_{F}^{2}=\left\|\mathsf{E}_{\mathbf{X}}\right\|_{F}^{2}+\sum_{j}\mathrm{Tr}\left(\boldsymbol{\Sigma}_{\mathbf{X}_{:j}}\right)=\left\|\mathsf{E}_{\mathbf{X}}\right\|_{F}^{2}+n_{2}\sum_{i}\frac{\mathbf{D}_{i}}{\alpha\mathbf{D}_{i}+\beta}.$
@@
(24) $\mathbb{E}\left\|\mathbf{X}\right\|_{*}=\frac{1}{2}\sum_{j}\mathbb{E}\left[\mathbf{X}_{:j}^{T}\omega({\mathsf{E}_{\mathbf{X}}})\mathbf{X}_{:j}\right]+\frac{1}{2}\|\mathsf{E}_{\mathbf{X}}\|_{*}.$
@@
(25) $\mathbb{E}\left\|\mathbf{X}\right\|_{*}=\left\|\mathsf{E}_{\mathbf{X}}\right\|_{*}+\frac{n_{2}}{2}\mathrm{Tr}\left(\left(\alpha{\mathbf{D}}+\beta\mathbf{I}\right)^{-1}\right)=\left\|\mathsf{E}_{\mathbf{X}}\right\|_{*}+\frac{n_{2}}{2}\sum_{i}\frac{1}{\alpha\mathbf{D}_{i}+\beta}.$

### v2

@@ 3.2.1 Prior distribution for $\mathcal{S}$
For the sparse component $\mathcal{S}$, we employ a Laplace prior distribution that induces $\ell_{1}$-norm regularization. This choice is motivated by the well-established connection between Laplace priors and sparsity promotion in the Bayesian framework [42]. Specifically, the prior density takes the form:
$p(\mathcal{S}|\theta_{2})\propto{\theta_{2}^{n}}\exp\left(-\theta_{2}\|\mathcal{S}\|_{1}\right),$ (10)
where $\theta_{2}>0$ is a scale parameter. The $\ell_{1}$-norm arises naturally as the convex envelope of the $\ell_{0}$ pseudo-norm, making it the tightest convex relaxation for sparse recovery problems. From a probabilistic perspective, this corresponds to assuming independent exponentially distributed entries in $\mathcal{S}$, which favors exact zeros in the MAP estimate while maintaining computational tractability through convex optimization.
Taking the logarithm of the prior distribution, we obtain:
$\log p(\mathcal{S}|\theta_{2})=-\theta_{2}\|\mathcal{S}\|_{1}+n\log\theta_{2}+C_{2},$ (11)
@@ 3.2.2 Prior distribution for $\mathcal{L}$
For the variable $\mathcal{L}$, we employ a particular Gibbs prior  [29] to promote a low-rank structure in $\mathcal{L}$. This prior takes the form of an exponential distribution with a tensor nuclear norm penalty, acting as a convex surrogate for the tensor average rank. It encourages $\mathcal{L}$ to have a low-rank representation, which is often suitable for capturing the underlying low-dimensionality in the data. The prior distribution is given by:
$p(\mathcal{L}|\theta_{3})\propto{\theta_{3}^{n}}\exp\left(-\theta_{3}\left\|\mathcal{L}\right\|_{*}\right).$ (12)
@@ 3.3 The hyper-prior distribution
In the field of statistical modeling, the Gamma distribution has obtained significant attention as a versatile prior distribution for hyperparameters, particularly in Bayesian frameworks [2, 3, 4, 38, 43]. The choice of a Gamma distribution as the prior for the hyperparameter $\theta_{i}$ is driven by two key reasons. First, it serves as a conjugate prior for precision parameters in exponential family distributions. For instance, when $\theta_{i}$ controls the precision of a Gaussian likelihood $p(x|\theta_{i})\sim\mathcal{N}(0,\theta_{i}^{-1})$, the Gamma prior ensures the posterior distribution remains a Gamma distribution. This conjugacy simplifies posterior calculations in Bayesian inference, enabling efficient automatic updates of hyperparameters. Second, $\theta_{i}$ typically represents positive physical quantities like precision or rate. The Gamma distribution’s support on $(0,+\infty)$ naturally aligns with this positivity constraint, eliminating the need for artificial non-negativity restrictions.
We assign independent Gamma priors to the hyperparameters $\theta_{i}$, which correspond to the mutually independent components $\mathcal{E}$, $\mathcal{S}$, and $\mathcal{L}$ in the model. This hierarchical structure preserves model consistency while enabling efficient computation. The independence assumption further facilitates automatic feature selection by factorizing the posterior distribution into marginal products over each $\theta_{i}$. Hence, we have
@@ 3.4 Joint distribution
We remark that, in the MAP framework, the hyperparameters $\boldsymbol{\theta}$ must be either pre-specified or estimated prior to the estimation of $\mathcal{L}$ and $\mathcal{S}$. For a more comprehensive estimation that includes the hyperparameters, the joint maximum a posteriori (JMAP) estimation is employed:
$(\mathcal{S}^{\dagger},\mathcal{L}^{\dagger},\boldsymbol{\theta}^{\dagger})=\arg\max_{\mathcal{S},\mathcal{L},\boldsymbol{\theta}}p(\mathcal{S},\mathcal{L},\boldsymbol{\theta}|\mathcal{X})=\mathop{\rm argmax}_{\mathcal{S},\mathcal{L},\boldsymbol{\theta}}\frac{p(\mathcal{X},\mathcal{S},\mathcal{L},\boldsymbol{\theta})}{p(\mathcal{X})}.$ (13)
For simplicity, we assume independence among the hyperparameters, allowing us to express the joint density function of the variables $\mathcal{X}$, $\mathcal{S}$, $\mathcal{L}$, and $\boldsymbol{\theta}$ as:
@@
In the literature [4, 2, 38, 3, 43], the Gamma distribution is commonly adopted as a prior for the hyperparameters $\theta_{i}$ ($i=1,2,3$) due to its conjugacy with certain likelihood functions, which facilitates analytical tractability in Bayesian inference. However, prior knowledge about the shape and scale parameters ($a_{\theta_{i}}$ and $b_{\theta_{i}}$) of the Gamma distribution is often lacking. To address this, a non-informative prior can be implemented by adopting an improper uniform prior distribution, defined as $p(x)\propto 1$ for $x\in\{\theta_{i}\mid i=1,2,3\}$ over the positive real line [43, 15]. Hence we have
$p(\mathcal{X},\mathcal{S},\mathcal{L},\boldsymbol{\theta})\propto\theta_{1}^{n/2}\theta_{2}^{n}\theta_{3}^{n}\exp\left(-\frac{\theta_{1}}{2}\left\|\mathcal{X}-\mathcal{S}-\mathcal{L}\right\|_{F}^{2}-\theta_{2}\left\|\mathcal{S}\right\|_{1}-\theta_{3}\left\|\mathcal{L}\right\|_{*}\right).$ (14)
@@ 4 Variational Bayesian inference
In Bayesian modeling, inference involves conditioning on observed data $\mathcal{X}$ and estimating the posterior density $p(\mathcal{S},\mathcal{L},\boldsymbol{\theta}|\mathcal{X})$. This task can be tackled via Markov Chain Monte Carlo (MCMC) sampling or optimization approaches. However, in this paper, we adopt variational inference as the methodological framework to approximate the latent variables $\mathcal{L}$ and $\mathcal{S}$, along with the parameter vector $\boldsymbol{\theta}$.
@@ 4.1 Kullback-Leibler divergence and evidence lower bound
The central goal of variational inference is to identify an optimal variational density $q(\mathcal{S},\mathcal{L},\boldsymbol{\theta})$ that closely approximates the posterior density $p(\mathcal{S},\mathcal{L},\boldsymbol{\theta}|\mathcal{X})$, thereby facilitating efficient inference on the latent variables and parameters [8].
Within this framework, we define a family of densities $\mathcal{Q}$ over the latent variables and parameters. Each candidate $q(\mathcal{S},\mathcal{L},\boldsymbol{\theta})\in\mathcal{Q}$ represents an approximation to the true posterior. The optimal candidate is chosen by minimizing the Kullback-Leibler (KL) divergence from the true posterior:
$\text{KL}(q(\mathcal{S},\mathcal{L},\boldsymbol{\theta})\|p(\mathcal{S},\mathcal{L},\boldsymbol{\theta}|\mathcal{X}))$ $=\int_{\mathcal{S},\mathcal{L},\boldsymbol{\theta}}q(\mathcal{S},\mathcal{L},\boldsymbol{\theta})\log\left(\frac{q(\mathcal{S},\mathcal{L},\boldsymbol{\theta})}{p(\mathcal{S},\mathcal{L},\boldsymbol{\theta}|\mathcal{X})}\right)\,d\mathcal{L}\,d\mathcal{S}\,d\boldsymbol{\theta}$ (15)
@@
According to (13), the posterior density $p(\mathcal{S},\mathcal{L},\boldsymbol{\theta}|\mathcal{X})$ is the ratio between $p(\mathcal{X},\mathcal{S},\mathcal{L},\boldsymbol{\theta})$ and $p(\mathcal{X})$. The density $p(\mathcal{X})$ involves integrating out the latent variables from the joint density. Unfortunately, this integration is often intractable, rendering direct computation of the posterior challenging. Expand the condition density, we have
@@
The second term is independent of latent variables and hyperparameters; therefore, it’s just a constant in the minimization problem, and we can ignore this term. To circumvent the intractability, we optimize an alternative objective that is equivalent to the KL divergence up to an additive constant. Specifically, we minimize the first term on the right-hand side of the equation, which constitutes the evidence lower bound (ELBO), denoted $\mathcal{J}(q(\mathcal{S},\mathcal{L},\boldsymbol{\theta}))$
$\mathcal{J}(q(\mathcal{S},\mathcal{L},\boldsymbol{\theta}))\equiv\mathbb{E}_{q(\mathcal{S},\mathcal{L},\boldsymbol{\theta})}\left[\log\left(\frac{p(\mathcal{X},\mathcal{S},\mathcal{L},\boldsymbol{\theta})}{q(\mathcal{S},\mathcal{L},\boldsymbol{\theta})}\right)\right].$ (16)
@@
$q^{\dagger}(\mathcal{S},\mathcal{L},\boldsymbol{\theta})=\mathop{\rm argmax}_{{q(\mathcal{S},\mathcal{L},\boldsymbol{\theta})\in\mathcal{Q}}}\mathcal{J}(q(\mathcal{S},\mathcal{L},\boldsymbol{\theta})).$ (17)
@@ 4.2 Mean-field variational family
The selection of variational densities $q(\mathcal{L})$, $q(\mathcal{S})$, and $q(\theta_{i})$ is importance. For $q(\mathcal{L})$ and $q(\mathcal{S})$, we adopt normal distributions due to their versatility and analytical convenience. The choice of the variational density $q(\theta_{i})$ as a Gamma distribution is motivated by the conjugacy properties derived from the likelihood function (Eq. 9) and the prior distributions specified in Eqs. (10) and (12). These indicate that the posterior distribution of $\theta_{1}$ and the conditional posteriors of $\theta_{i}$ for $i=2,3$ follow Gamma distributions. Since the Gamma distribution is conjugate to itself, selecting $q(\theta_{i})$ as a Gamma density ensures compatibility with the posterior, facilitating efficient variational inference.
Let $\mathcal{Q}_{\mathcal{G}}$ denote the set of Gamma densities for the hyperparameters $\theta_{i}$ ($i=1,2,3$), and $\mathcal{Q}_{\mathcal{N}}$ denote the set of multivariate normal densities over the tensor space $\mathbb{R}^{n_{1}\times n_{2}\times n_{3}}$. The overall variational family $\mathcal{Q}$ can be expressed as the Cartesian product of these sets: $\mathcal{Q}=\mathcal{Q}_{\mathcal{N}}\times\mathcal{Q}_{\mathcal{N}}\times\mathcal{Q}_{\mathcal{G}}$.
@@ 4.3 Laplacian approximation
In (14), the non-quadratic properties inherent in both the $\ell_{1}$ norm of $\mathcal{S}$, which represents the sum of the absolute values of all elements, and the tensor nuclear norm of $\mathcal{L}$, which is the weighted sum of its singular values, pose significant obstacles for direct optimization within standard density families. These non-quadraticities complicate direct inference procedures, rendering them computationally intractable. To address this, we utilize the Laplace approximation method, involving mean calculation, variance estimation, and density function construction, to approximate the density with a Gaussian distribution form.
@@
where $\mathcal{Q}_{\mathcal{N}}$ is the set of all the density functions for the Gaussian distribution. According to Gibbs’ inequality, for any two probability distributions $q(x)$ and $p(x)$ over a domain $\Omega$, the following holds:
@@ 4.3.1 Absolute value function
In the paper, we utilize the sans serif font $\mathsf{E}$ accompanied by a subscripted variable $x$ to denote the expectation of a random variable $x$. The solution is given by:
@@
$q(x)=\mathcal{N}\left(\mathsf{E}_{x},\frac{|\mathsf{E}_{x}|}{|\mathsf{E}_{x}|+\beta}\right).$ (18)
@@
$\mathsf{E}|x|=|\mathsf{E}_{x}|+\frac{1}{2(|\mathsf{E}_{x}|+\beta)}.$ (19)
@@ 4.3.2 Nuclear norm
$\mathsf{E}_{\mathbf{X}}=\arg\min_{\mathbf{X}}\left(\frac{\alpha}{2}\left\|\mathbf{X}-\mathbf{A}\right\|_{F}^{2}+\beta\left\|\mathbf{X}\right\|_{*}\right).$
@@
$\left\|\mathbf{X}\right\|_{*}\approx\frac{1}{2}\text{Tr}(\omega(\mathsf{E}_{\mathbf{X}})\mathbf{X}\mathbf{X}^{T})+\frac{1}{2}\|\mathsf{E}_{\mathbf{X}}\|_{*}.$ (20)
@@
$\boldsymbol{\Sigma}_{\mathbf{X}_{:j}}={\mathbf{U}}{\mathbf{D}}\left(\alpha{\mathbf{D}}+\beta\mathbf{I}\right)^{-1}{\mathbf{U}}^{T}.$ (21)
@@
$q(\mathbf{X})=\prod_{j}\mathcal{N}(\mathbf{X}_{:j}|\mathsf{E}_{\mathbf{X}_{:j}},{\mathbf{U}}{\mathbf{D}}\left(\alpha{\mathbf{D}}+\beta\mathbf{I}\right)^{-1}{\mathbf{U}}^{T}).$ (22)
@@
$\mathbb{E}\left\|\mathbf{X}\right\|_{F}^{2}=\left\|\mathsf{E}_{\mathbf{X}}\right\|_{F}^{2}+\sum_{j}\mathrm{Tr}\left(\boldsymbol{\Sigma}_{\mathbf{X}_{:j}}\right)=\left\|\mathsf{E}_{\mathbf{X}}\right\|_{F}^{2}+n_{2}\sum_{i}\frac{\mathbf{D}_{i}}{\alpha\mathbf{D}_{i}+\beta}.$ (23)
@@
$\mathbb{E}\left\|\mathbf{X}\right\|_{*}=\frac{1}{2}\sum_{j}\mathbb{E}\left[\mathbf{X}_{:j}^{T}\omega({\mathsf{E}_{\mathbf{X}}})\mathbf{X}_{:j}\right]+\frac{1}{2}\|\mathsf{E}_{\mathbf{X}}\|_{*}.$ (24)
@@
$\mathbb{E}\left\|\mathbf{X}\right\|_{*}=\left\|\mathsf{E}_{\mathbf{X}}\right\|_{*}+\frac{n_{2}}{2}\mathrm{Tr}\left(\left(\alpha{\mathbf{D}}+\beta\mathbf{I}\right)^{-1}\right)=\left\|\mathsf{E}_{\mathbf{X}}\right\|_{*}+\frac{n_{2}}{2}\sum_{i}\frac{1}{\alpha\mathbf{D}_{i}+\beta}.$ (25)